OpenStudy (campayne):

Which of the following is an odd function? f(x) = x3 + 5x2 + x f(x) = square root of x f(x) = x2 + x f(x) = –x

OpenStudy (campayne):

it should be noted that i have only five minutes as of right now to complete this test

OpenStudy (campayne):

HERES ANOTHR ONE IF YOU WOULD BE KIND PLEASE If f(x)= 1/9 x - 2 , what is f–1(x)? f–1(x) = 9x + 18 f-1(x) = 1/9x + 2 f–1(x) = 9x + 2 f-1(x) = -2x + 1/9

OpenStudy (campayne):

TWO MIN GUYS PLEASE

OpenStudy (campayne):

MIN AND HALF

OpenStudy (campayne):

SOMEBODY LOOKED BUT LEFT NOO PLEASSEEEE

OpenStudy (campayne):

ONE MIN

OpenStudy (campayne):

NOOOOOOOOOOO

OpenStudy (campayne):

THIRTY SECONDS

OpenStudy (campayne):

TEN SECONDS

OpenStudy (campayne):

oh well i got a 72% good enough for me

OpenStudy (triciaal):

Just saw this but anyway, we cannot help you on a test. considered cheating. against the code of conduct.